Given: x = 1 - cost and y = 1 - sint. Find y’.


1
Expert's answer
2022-02-24T11:47:48-0500

It follows that

"y'_x=\frac{y'_t}{x'_t}=\frac{-\cos t}{\sin t}=-\cot t."